OpenStudy (anonymous):

may someone please explain to me what an open equation is i completly forgot

OpenStudy (anonymous):

Open equation is when its an equation where you can get a clear variable answer. @Thixchix$3 fan and medal

Nnesha (nnesha):

http://openstudy.com/study#/updates/52371f3de4b0af32a079babf this might be helpfl

OpenStudy (anonymous):

is (1,5) a solution if the equation y=x-2?

OpenStudy (anonymous):

please give an explanation also..And how do i give medals?

Nnesha (nnesha):

okay just replace x and y by their values in order pair first value is x and 2nd one is y like this ( x , y) so substitute x and y values into y = x-2 if you get both side equal then yes

Nnesha (nnesha):

http://goo.gl/TJE4k3 medal :) *click on best response*

OpenStudy (anonymous):

i already did

Nnesha (nnesha):

hah thanks that was answer of your question

Nnesha (nnesha):

okay back to math question

OpenStudy (anonymous):

i have one more..sorry

Nnesha (nnesha):

its okay :)

Nnesha (nnesha):

what u got for last one ?

OpenStudy (anonymous):

which o the folllowing equation has 3 as a solution?

OpenStudy (anonymous):

-2b=6 4a-5=12 5y=y+8 z-15=-2

OpenStudy (anonymous):

those are the options

Nnesha (nnesha):

there is 2 way to solve this u can either replace variables by 3 and if yo u get both side equal then that is your answer or you can solve all equations for variables which one is easy for you ?

OpenStudy (anonymous):

replace variables by 3

Nnesha (nnesha):

okay so replace variables by 3 if you get both side equal that equation have 3 as a soultion

Nnesha (nnesha):

got it ?? let me know if you have aaa question ?

OpenStudy (anonymous):

the answer is 5y=y+8

Nnesha (nnesha):

 replace both y by 3

OpenStudy (anonymous):

so what is the answer??

Nnesha (nnesha):

 what is 5 times 3 = ?? and 3 + 8 = ??